Please help me solve this problem that I am having trouble with:
2/5x=-10x

Answer by malakumar_kos@yahoo.com(315)


2/5x=-10x solve
50x^2 = 2

x^2 = 2/50
= 1/25
x^2 = 1/25
x = +-1/5 (taking square root on both sides)